# Boundary crisis and long transients of the Atlantic overturning circulation mediated by an edge state

### v1

@@ III Beyond stable states: A global view
Here may depend explicitly on time via the -dimensional external forcing input [44]. For fixed external forcing , the dynamics is given by the so-called frozen system [44]. Multistability is characterised by the coexistence of multiple attractors for given . In our case, with ppm, there are two chaotic attractors corresponding to the ON and OFF states. Each attractor possesses a basin of attraction, i.e., a set of initial conditions that evolve towards it as . Since PlaSim-LSG is fully deterministic, the asymptotic state is thus uniquely determined by the initial condition in the absence of perturbations333More generally, stochastic parameterisations of unresolved processes change this picture and require a stochastic description.
The two basins of attraction must be separated by a basin boundary of codimension with respect to the dimension (number of degrees of freedom) of the system. The basin boundary between the two AMOC states in PlaSim-LSG is thus a high-dimensional set in the model phase space. As was shown for conceptual [48] and intermediate-complexity [10] climate models, the basin boundary can be fractal with almost full phase space dimension (). The basin boundary is crucial in the context of critical transitions because it marks the threshold in phase space where the dominant feedback changes from a stabilizing (negative) feedback that exerts a restoring force towards the original attractor to a destabilizing (positive) feedback that drives a self-perpetuating transition to a competing attractor.
Even if the “curse of dimensionality” [91] currently prevents computing the quasipotential of high-dimensional systems, we can learn much about its structure by analysing edge states as important regions therein. Since edge states are unstable, they generally cannot be found by direct simulation. In the following, we describe how edge states can be nonetheless computed purely based on forward integration of the model.
@@ III.1 Finding the basin boundary
How can we locate the basin boundary between the two competing AMOC states in a high-dimensional climate model? We need a pair of initial conditions and that converge to the ON and OFF state, respectively. Interpolating along a straight line in phase space between and necessarily leads to crossing a segment of the basin boundary lying somewhere in between – provided that the interpolation is performed in the full phase space including all degrees of freedom (prognostic variables).
Here, we take two model restart files from previous PlaSim-LSG simulations [79] as initial conditions and . The simulations were performed at CO2 with differing vertical diffusivity profiles, which strongly affected the AMOC strength [79]. In our model configuration (at , see Supplemental Information for details), these two initial conditions are located near the ON and OFF state, respectively, and converge to these.
Now, interpolating along all variables with in the model restart files (), we compute new initial conditions with . Computing the meridional streamfunctions (Eq. (1)) for these initial conditions shows that the states are monotonically decreasing in AMOC strength with increasing (Fig. 5a). The differences in AMOC strength between adjacent states are not equidistant, reflecting that the AMOC strength is a nonlinear mapping of the full phase space.
From the initial conditions , we run parallel simulations for 200 years each. While the trajectory initialised at remains close to the ON state in AMOC strength, all other trajectories converge to the OFF state. This implies that a part of the basin boundary is located between and in phase space. This pair of initial conditions constitutes the starting points of our edge tracking procedure.
@@ III.2 Edge tracking algorithm
Bisection. Between two initial conditions converging to attractors and , respectively, bisect repeatedly along a straight line in phase space to obtain two new initial conditions that are less than a distance apart while still converging to different attractors (one to and the other to ).
@@
Here the distance measures could be the Euclidean distance in a normalised phase space or any other appropriate measure of separation between the two states. We simply measure the difference in 10-year smoothed AMOC strength.
Running the algorithm yields two strings of trajectory segments that shadow the basin boundary on either side of it. By concatenating each of these strings in time and taking their average at each time point, we obtain a pseudotrajectory that approximates a trajectory on the basin boundary. The repeated rebisection of initial conditions thereby counteracts the instability that causes any trajectory initialised near the boundary to eventually diverge from it. Based on the property of edge states as attracting sets when restricted to the boundary, the pseudotrajectory is expected to converge to an edge state. If this state supports chaos, the algorithm converges to this chaotic set and successively populates its invariant measure.
@@ III.3 Converging to an edge state
For the first few iterations of edge tracking, spanning about 200 years, the resulting pseudotrajectory (hereafter called edge trajectory) decreases in AMOC strength from to about . Subsequently, the edge trajectory begins a series of large AMOC oscillations (Fig. 5c). The quasiperiodic oscillations vary in amplitude from 3 to , with a mean period of years (estimated from 10 peaks). This behaviour persists until the edge tracking was stopped after around 1400 years (39 iterations).
@@
III.4 Reduced phase space
Looking at the one-dimensional AMOC timeseries (Fig. 5c) gives the impression that the edge trajectory oscillates back and forth between the ON and OFF states. However, visualizing the dynamics in a reduced phase space clarifies that the edge state is separated from the attractors (Fig. 6).
Determining a suitable low-dimensional projection of the -dimensional dynamics is challenging due to the countless possible combinations of variables. Based on our physical understanding of the AMOC, we consider the zonally averaged salinity field in the Atlantic. An empirical orthogonal function (EOF) analysis [92] combining of edge tracking simulations shows that this field contains sufficient information to disentangle the dynamics (Figs. S4 and S5 of the Supplemental Information). Specifically, the two leading EOFs reveal that most of the variance is explained by a meridional salinity dipole in the upper and a vertical dipole in the North Atlantic. From this we derive a reduced phase space spanned by the following three variables:
The meridional salinity gradient (SG) in the Atlantic, measured as the mean salinity difference between 0-20 N and 40-80 N in the top (omitting the top ),
The vertical SG in the North Atlantic, defined as the mean salinity difference between the depths 100- and 1000- at 46-66∘N,
The deep North Atlantic salinity anomaly, defined as the mean salinity anomaly relative to in the Atlantic basin north of and below depth.
@@ III.4 Reduced phase space
Viewing the trajectories of Fig. 5 in the reduced phase space, we see that each of the ON, OFF, and edge states occupies a distinct region (Fig. 6). The edge state has a higher vertical SG and fresher deep North Atlantic than both the ON and OFF states. The OFF state has the saltiest deep North Atlantic and largest meridional SG. While the ON state covers a relatively small volume of the reduced phase space, the AMOC oscillations of the edge state are clearly seen as loops in the meridional-vertical SG plane. Also the OFF state exhibits relatively large internal variability that is captured in this projection but not in the AMOC strength. This low-frequency variability on multi-centennial timescales is caused by global inter-basin salt exchanges (not shown).
@@ IV.1 Energetics
The first question is whether the edge state energetically fulfills steady state conditions, requiring an approximately vanishing global energy budget for the coupled climate system and its subcomponents [93, 94]. Indeed, both the radiative balance at the top of the atmosphere as well as the globally integrated net surface heat flux between the ocean and atmosphere are close to zero on the edge state (comparable to the ON and OFF states; Fig. 7a).
The meridional heat transport of the ocean and atmosphere combined is nearly identical for the ON, OFF and edge states, despite differences in the ocean circulation (7b). This means that the atmosphere largely compensates for changes in oceanic heat transport [95], manifesting the Bjerknes compensation [96, 97] also reported in previous studies on the AMOC variability and collapse [98, 86]. Because of the AMOC, the Atlantic Ocean is the only ocean basin with a northward oceanic heat transport on both hemispheres, causing an asymmetry of the oceanic meridional heat transport. A reduced AMOC thus decreases this asymmetry, as we observe for the OFF state (Fig. 7c, lower panel). Interestingly, the change in the atmospheric transport slightly overcompensates the reduction in the oceanic transport (Fig. 7c, upper panel). The time-averaged oceanic meridional heat transport of the edge state lies in between that of the ON and OFF states, though the AMOC oscillations cause temporal variations of more than 0.1 PW especially in the northern mid-latitudes (Fig. 7d).
@@
as an approximate energy measure to compare the oceanic specific potential energy among the different AMOC states. Here m is the maximum depth of the sea floor, is the depth coordinate (positive downwards), and is the horizontally integrated density across the ocean ( below the sea floor).
The edge state has a significantly higher centre of mass – and thus specific potential energy – compared to the two attractors. This aligns with the situation in a global ocean model, where the dynamic enthalpy of the edge state was shown to be elevated [57]. Here, the ON and OFF states have a comparable centre of mass, with the OFF state exhibiting multi-centennial variability in due to global salt exchanges, as also observed in Fig. 6.
@@
IV.2 Eccentric observables
Since the edge state lies on the basin boundary between the ON and OFF states, one might expect that its climate lies somewhere in between that of the ON and OFF states, too. For example, the AMOC strength and meridional heat transport on the edge state oscillate between the strong and weak AMOC states. At the same time, the edge state features a fresher deep North Atlantic than both attractors, and a higher centre of mass. In a high-dimensional system like our climate model, there may be many directions – which we term eccentric observables – in which the edge state lies “outside” of the ON and OFF states. These directions in phase space could be particularly relevant for detecting EWS, as proposed in Ref. [99] and discussed in sec. IV.2. Along transition paths (provided that they pass via the vicinity of the edge state), we expect eccentric observables to undergo non-monotonic excursions in these observables. Thus, the signal of a transition could initially have the opposite sign of the anticipated change.
In most ocean regions, the time-averaged sea surface salinity (SSS) and sea surface temperature (SST) of the edge state lie in between that of the ON and OFF states (Fig. 8). However, almost the entire Arctic Ocean is saltier and denser in the upper ocean relative to the attractors. Parts of the NorS are warmer than both attractors, and the northwestern Pacific Ocean is warmer and saltier on the edge state. Other eccentric observables include the sea ice cover in the IrmS and the surface freshwater flux in the NorS.
@@ IV.3 Drivers of the unstable oscillations
The most prominent dynamical feature of the edge state are the large AMOC oscillations with a period of around 120 years and an amplitude of up to between 46-66∘N (Fig. 5c). At 26∘N, the AMOC oscillations are qualitatively similar but have a smaller amplitude (Fig. 9a). Together with the overturning strength, many other climate observables oscillate at this frequency (Fig. 9). What drives the unstable oscillations?
@@
Key processes known to interact with the AMOC are North Atlantic deep convection, sea ice formation, and high-latitude surface freshwater fluxes. As shown in Fig. 3, the transition from ON to OFF in our model is characterised by a shutdown of deep convection sites in the Labrador and NorSs. On the edge state, deep convection persists in the NorS, though the mean convection depth oscillates slightly with the AMOC strength (Fig. 9b). Convection is inactive in the IrmS. In the LabS, deep convection undergoes large variations, switching on and off in close correspondence with the AMOC strength.
@@
To systematically investigate how the various oscillating variables are related to each other in time, we compute lag correlations between the AMOC strength at 46-66∘N and all other variables, considering time lags between and 120 years. We select variables whose 3-year smoothed timeseries has a maximum lag correlation above 0.8 in absolute value. For these variables, we compute correlation values also for the unfiltered timeseries (annual resolution), giving the values reported in the following and in Fig. 10.
@@
The sea ice cover fraction oscillates strongly in the LabS and, with a phase shift of around , in the IrmS (Fig. 9c). In both regions, the sea ice retreats almost entirely during the AMOC maximum. In the NorS, there is little sea ice on the edge state at all times. A clear phase shift is also seen between the mean upper ocean salinity (top 1000 m) of the LabS and NorS (Fig. 9d). Generally, the salinity changes could be caused by horizontal advection, convection, or surface freshwater fluxes. We find a strong southward freshwater transport through the Denmark Strait between Greenland and Iceland, oscillating in anti-phase with the AMOC (Fig. 9e), as well as a large amplitude in precipitation minus evaporation (PE) over the LabS (Fig. 9f). Furthermore, the magnitude of the wind-driven southward Ekman transport in the North Atlantic (50-60∘N) is negatively correlated with the AMOC strength at a lead time of four years (Figs. 9h and 10). Around the AMOC minimum on the edge state, the Ekman transport is stronger compared to both the ON and OFF states, indicating a potential role of the wind stress in triggering AMOC weakening by counteracting the northward overturning current [78] (see Fig. S2 of the Supplemental Information). In the atmosphere, temperature, precipitation and winds likewise display variability on the 120-year timescale (Fig. 9g-i), though correlations with the AMOC strength are lower due to the much higher interannual variability in the atmosphere compared to the ocean.
A key observation is that the upper ocean salinity, deep convection, sea ice, and PE in the LabS all lead the AMOC by 6-8 years, with correlations ranging between 0.78 and 0.92 (in absolute value). PE averaged across the entire North Atlantic (between 50-80∘N) has an even larger lead time of 11 years, though the correlation is less strong. The AMOC strength measured at 26∘N follows the AMOC at 46-66∘N by 6 years, and the overturning component of the freshwater export at the Atlantic southern border (34∘S) [42] lags behind by 26 years.
@@
Our analysis thus shows that the LabS is a key region in driving the AMOC oscillations on the edge state. The fact that the salt and volume transport in the tropical Atlantic lags the AMOC strength further north suggests that the salt-advection feedback does not initiate the oscillations, though it likely plays an important role in amplifying them. Rather, ocean-cryosphere-atmosphere interactions in the North Atlantic appear crucial for triggering the AMOC cycles.
@@
To gain further insight into the processes at play, we now divide each AMOC cycle into four phases (A–minimum, B–rise, C–maximum, D–decline; see Fig. 11) and consider time averages for each phase over the final five oscillations of the edge trajectory. In phase A, the LabS is ice-covered, preventing deep convection and thus maintaining a weak AMOC. However, deep water formation in the NorS ensures that the AMOC is not as weak as in the OFF state. In phase B, sea ice retreats in the LabS, which allows the ocean to release heat to the atmosphere and consequently deep convection to be activated. The salt-advection feedback kicks in, supplying warm and salty water to the LabS, enhancing sea ice retreat and convection up to the AMOC maximum in phase C. Then, however, the salinity and convection in the LabS start to decrease, along with sea ice expansion. One possible reason for this reversal could be the strong freshwater influx from the Arctic Ocean through the Denmark Strait, which peaks in phase C and reaches to the LabS. The freshwater flow is concentrated in the upper ocean, implying that it can disrupt convection by freshening the upper water column. Another explanation could involve surface fluxes of heat and freshwater. In phase D, sea ice rapidly expands to cover the entire LabS and convection shuts down, causing the AMOC decline.
@@
Fully deciphering the oscillation mechanism goes beyond the scope of this study. Nonetheless, we can identify multiple competing processes that could produce cyclic behaviour: a competition between sea ice and convection in the LabS, a competition between salt advection by the AMOC and freshwater advection from the Arctic Ocean, as well as a competition between deep water formation sites in the Labrador and NorSs. The competition between deep convection sites could explain the anti-phase pattern observed in precipitation between Greenland and the United Kingdom (Fig. 9g) as well as in surface air temperatures between the Greenland-Iceland-Norwegian (GIN) Seas and the rest of the high northern latitudes (Fig. 11c). During phase A, at the AMOC minimum, air temperatures are warmer over the GIN Seas than during phase C, in contrast with the surrounding areas, especially the LabS. We hypothesise that this is because the NorS is the only deep convection zone in the North Atlantic during phase A, directing the meridional heat transport to that region and reducing cold inflows from the Arctic Ocean.
@@
Exploring the dynamics on the edge state thus reveals distinct modes of climate variability that are absent in the ON and OFF attractors. The edge tracking method allowed to capture the centennial climate oscillations even though they are asymptotically unstable. Despite being unstable, such oscillations can persist for hundreds of years on Earth system trajectories evolving near the edge state, as we show in the following.

### v2

@@ III Beyond stable states: A global view
Here may depend explicitly on time via the -dimensional external forcing input [44]. For fixed external forcing , the dynamics is given by the so-called frozen system [44]. Multistability is characterised by the coexistence of multiple attractors for given . In our case, with ppm, there are two chaotic attractors corresponding to the ON and OFF states. Each attractor possesses a basin of attraction, i.e., a set of initial conditions that evolve towards it as . Since PlaSim-LSG is fully deterministic, the asymptotic state is thus uniquely determined by the initial condition in the absence of perturbations222More generally, stochastic parameterisations of unresolved processes change this picture and require a stochastic description.
The two basins of attraction must be separated by a basin boundary of dimension with respect to the dimension (number of degrees of freedom) of the system. The basin boundary between the two AMOC states in PlaSim-LSG is thus a high-dimensional set in the model state space. As was shown for conceptual [48] and intermediate-complexity [10] climate models, the basin boundary can be fractal with almost full state space dimension (). The basin boundary is crucial in the context of critical transitions because it marks the threshold in state space where the dominant feedback changes from a stabilizing (negative) feedback that exerts a restoring force towards the original attractor to a destabilizing (positive) feedback that drives a self-perpetuating transition to a competing attractor.
Even if the “curse of dimensionality” [91] currently prevents computing the quasipotential of high-dimensional systems, we can learn much about its structure by analysing edge states as important landmarks therein. Since edge states are unstable, they generally cannot be found by direct simulation or basic continuation. However, since they are unstable in only one direction (transversal to the basin boundary), edge states can nonetheless be computed purely based on forward integration of the model, as described in the following.
@@ III.1 Finding the basin boundary
How can we locate the basin boundary between two competing AMOC states in a high-dimensional climate model? We need a pair of initial conditions and that are attracted by the ON and OFF state, respectively. Interpolating along a straight line in state space between and necessarily leads to crossing a segment of the basin boundary lying somewhere in between – provided that the interpolation is performed in the full state space including all degrees of freedom (prognostic variables).
Here, we take two model restart files from previous PlaSim-LSG simulations [80] as initial conditions and . The simulations were performed at 285 ppm CO2 with differing vertical diffusivity profiles, which strongly affected the AMOC strength [80]. In our model configuration (at 360 ppm, see Supplemental Information for details), these two initial conditions are located near the ON and OFF state, respectively, and evolve to these.
Now, interpolating along all variables with in the model restart files (), we compute new initial conditions for . Computing the meridional streamfunctions (Eq. (1)) for these initial conditions shows that the states are monotonically decreasing in AMOC strength with increasing (Fig. 5a). The differences in AMOC strength between adjacent states are not equidistant, reflecting that the AMOC strength is a nonlinear mapping of the full state space.
From the initial conditions , we run parallel simulations for 200 years each. While the trajectory initialised at remains close to the ON state in AMOC strength, all other trajectories lead to the OFF state. This implies that a part of the basin boundary is located between and in state space. This pair of initial conditions constitutes the starting points of our edge tracking procedure.
@@ III.2 Edge tracking algorithm
Bisection. Between two initial conditions converging to attractors and , respectively, bisect repeatedly along a straight line in state space to obtain two new initial conditions that are less than a distance apart while still converging to different attractors (one to and the other to ).
@@
Here the distance measures could be the Euclidean distance in a normalised state space or any other appropriate measure of separation between the two states. We simply measure the difference in 10-year smoothed AMOC strength.
Running the algorithm yields two series of trajectory segments that shadow the basin boundary on either side of it. By concatenating the segments and averaging over both series at each time point, we obtain a pseudotrajectory that approximates a trajectory on the basin boundary. The repeated rebisection of initial conditions thereby counteracts the instability that causes any trajectory initialised near the boundary to eventually diverge from it. Based on the property of edge states as attracting sets when restricted to the boundary, the pseudotrajectory is expected to converge to an edge state. If this state supports chaos, the algorithm converges to this chaotic set and successively populates its invariant measure.
@@ III.3 Converging to an edge state
For the first few iterations of edge tracking, spanning about 200 years, the resulting pseudotrajectory (hereafter called edge trajectory) decreases in AMOC strength from 14 Sv to about 5 Sv. Subsequently, the edge trajectory begins a series of large AMOC oscillations (Fig. 5c). The quasiperiodic oscillations vary in amplitude from 3 to 10 Sv, with a mean period of years (estimated from 10 peaks). This behaviour persists until the edge tracking was stopped after around 1400 years (39 iterations).
@@
III.4 Reduced state space
Looking at the one-dimensional AMOC timeseries (Fig. 5c) gives the impression that the edge trajectory oscillates back and forth between the ON and OFF states. However, visualizing the dynamics in a reduced state space clarifies that the edge state is separated from the attractors (Fig. 6).
Determining a suitable low-dimensional projection of the -dimensional dynamics is challenging due to the countless possible combinations of variables. Based on our physical understanding of the AMOC, we consider the zonally averaged salinity field in the Atlantic. An empirical orthogonal function (EOF) analysis [92] combining 20 000 years of edge tracking simulations shows that this field contains sufficient information to disentangle the dynamics (Figs. S4 and S5 of the Supplemental Information). Specifically, the two leading EOFs reveal that most of the variance is explained by a meridional salinity dipole in the upper 1000 m and a vertical dipole in the North Atlantic. From this we derive a reduced state space spanned by the following three variables:
The meridional salinity gradient (SG) in the Atlantic, measured as the mean salinity difference between 0-20 N and 40-80 N in the top 1000 m (omitting the top 100 m),
The vertical SG in the North Atlantic, defined as the mean salinity difference between the depths 100-1000 m and 1000-3000 m at 46-66∘N,
The deep North Atlantic salinity anomaly, defined as the mean salinity anomaly relative to 35 g kg-1 in the Atlantic basin north of 50∘N and below 1000 m depth.
@@ III.4 Reduced state space
Viewing the trajectories of Fig. 5 in the reduced state space, we see that each of the ON, OFF, and edge states occupies a distinct region (Fig. 6). The edge state has a higher vertical SG and fresher deep North Atlantic than both the ON and OFF states. The OFF state has the saltiest deep North Atlantic and largest meridional SG. While the ON state covers a relatively small volume of the reduced state space, the AMOC oscillations of the edge state are clearly seen as loops in the meridional-vertical SG plane. Also the OFF state exhibits relatively large internal variability that is captured in this projection but not in the AMOC strength. This low-frequency variability on multi-centennial timescales is caused by global inter-basin salt exchanges (not shown).
@@ IV.1 Energetics
The first question is whether the edge state energetically fulfills steady state conditions, requiring an approximately vanishing global energy budget for the coupled climate system and its subcomponents [93, 94]. Indeed, both the radiative balance at the top of the atmosphere as well as the globally integrated net surface heat flux between the ocean and atmosphere are close to zero (comparable to the ON and OFF states; Fig. 7a).
The meridional heat transport of the ocean and atmosphere combined is nearly identical for the ON, OFF and edge states, despite differences in the ocean circulation (7b). This means that the atmosphere largely compensates for changes in oceanic heat transport [95], manifesting the Bjerknes compensation [96, 97] also reported in previous studies on the AMOC variability and collapse [98, 87]. Because of the AMOC, the Atlantic Ocean is the only ocean basin with a northward oceanic heat transport on both hemispheres, causing an asymmetry of the oceanic meridional heat transport. A reduced AMOC thus decreases this asymmetry, as we observe for the OFF state (Fig. 7c, lower panel). Interestingly, the change in the atmospheric transport slightly overcompensates the reduction in the oceanic transport (Fig. 7c, upper panel). The time-averaged Atlantic meridional heat transport of the edge state lies in between that of the ON and OFF states, though the AMOC oscillations cause temporal variations of more than 0.1 PW especially in the northern mid-latitudes (Fig. 7d).
@@
as an approximate energy measure to compare the oceanic specific potential energy among the different AMOC states [99]. Here m is the maximum depth of the sea floor, is the depth coordinate (positive downwards), and is the horizontally integrated density across the ocean ( below the sea floor).
The edge state has a significantly higher centre of mass – and thus specific potential energy – compared to the two attractors. This aligns with the situation in a global ocean model, where the dynamic enthalpy of the edge state was shown to be elevated [57]. In our case, the ON and OFF states have a comparable centre of mass, with the OFF state exhibiting multi-centennial variability in due to global salt exchanges, as also observed in Fig. 6.
@@
IV.2 Excursive observables
Since the edge state lies on the basin boundary between the ON and OFF states, one might expect that its climate lies somewhere in between that of the ON and OFF states, too. For example, in terms of AMOC strength and Atlantic meridional heat transport the edge state oscillates between the ON and OFF states. At the same time, the edge state features a fresher deep North Atlantic than both attractors, and a higher centre of mass. In a high-dimensional system like our climate model, there may be many directions – which we term excursive observables – in which the edge state lies outside of the interval bounded by the two attractors. These directions in state space could be particularly relevant for detecting EWS [100] and for evaluating transition probabilities via rare event techniques relying on a score function [101]. Along transition paths (provided that they pass via the vicinity of the edge state), we expect excursive observables to undergo non-monotonic excursions. Thus, the signal of a transition could initially have the opposite sign of the anticipated change.
In most ocean regions, the time-averaged sea surface salinity (SSS) and sea surface temperature (SST) of the edge state lie in between that of ON and OFF states (Fig. 8). However, almost the entire Arctic Ocean is saltier and denser in the upper ocean relative to the attractors. Parts of the NorS are warmer than both attractors, and the northwestern Pacific Ocean is warmer and saltier on the edge state. Other excursive observables include the sea ice cover in the IrmS and the surface freshwater flux in the NorS.
@@ IV.3 Drivers of the unstable oscillations
The most prominent dynamical feature of the edge state are the large AMOC oscillations with a period of around 120 years and an amplitude of up to 10 Sv between 46-66∘N (Fig. 5c). At 26∘N, the AMOC oscillations are qualitatively similar but have a smaller amplitude (Fig. 9a). Together with the overturning strength, many other climate observables oscillate at this frequency (Fig. 9). What drives the unstable oscillations?
@@
In PlaSim-LSG, the transition from ON to OFF is characterised by a shutdown of all deep convection sites in the LabS, IrmS and NorS (Fig. 3). On the edge state, deep convection persists in the NorS, with some variation linked to the AMOC (Fig. 9b). In the LabS, deep convection undergoes large oscillations, switching on and off in close correspondence with the AMOC strength. Convection is inactive in the IrmS.
@@
To relate different oscillating variables in time, we compute lag correlations between the AMOC strength at 46-66∘N and other variables, considering time lags between and 120 years. We select variables whose 3-year smoothed timeseries has a maximum lag correlation above 0.8 in absolute value. For these variables, we compute correlation values also for the unfiltered timeseries (annual resolution), giving the values reported in the following and in Fig. S2 (Supplemental Information).
@@
The sea ice cover fraction oscillates strongly in the LabS and, with a phase shift of around , in the IrmS (Fig. 9c). In both regions, the sea ice retreats almost entirely during the respective minimum. In the NorS, there is little sea ice on the edge state at all times. A clear phase shift is also seen between the mean upper ocean salinity (top 1000 m) of the LabS and NorS (Fig. 9d). Generally, the salinity changes could be caused by horizontal advection, convection, or surface freshwater fluxes. We find a strong southward freshwater transport through the Denmark Strait between Greenland and Iceland, oscillating in anti-phase with the AMOC (Fig. 9e), as well as a large amplitude in precipitation minus evaporation (PE) over the LabS (Fig. 9f). Furthermore, the magnitude of the wind-driven southward Ekman transport in the North Atlantic (50-60∘N) is negatively correlated with the AMOC strength at a lead time of four years (Figs. 9h and S2, Supplemental Information). Around the AMOC minimum on the edge state, the Ekman transport is stronger compared to both the ON and OFF states, indicating a potential role of the wind stress in triggering an overturning decline [79] (see Fig. S6 of the Supplemental Information). In the atmosphere, temperature, precipitation and winds likewise display variability on the 120-year timescale (Fig. 9g-i), though correlations with the AMOC strength are lower due to the much higher interannual variability in the atmosphere compared to the ocean.
@@
A key observation is that the upper ocean salinity, deep convection, sea ice, and PE in the LabS all lead the AMOC by 6-8 years, with lag correlations ranging between 0.78 and 0.92 (in absolute value, see Fig. S2, Supplemental Information). PE averaged across the entire North Atlantic (between 50-80∘N) has an even larger lead time of 11 years, though the correlation is less strong. The AMOC strength measured at 26∘N follows the AMOC at 46-66∘N by 6 years, and the overturning component of the freshwater export at the Atlantic southern border (34∘S) [42] lags behind by 26 years. Our analysis thus shows that the LabS is a key region in driving the AMOC oscillations on the edge state. The fact that the salt and volume transport in the tropical Atlantic lags the AMOC strength further north suggests that the salt-advection feedback does not initiate the oscillations, though it likely plays an important role in amplifying them. Rather, ocean-ice-atmosphere interactions in the North Atlantic and Arctic appear crucial for triggering the AMOC cycles.
@@
To gain further process understanding, we now divide each AMOC cycle into four phases (A–minimum, B–rise, C–maximum, D–decline; see Fig. 10) and consider time averages for each phase over the final five oscillations of the edge trajectory. In phase A, the LabS is ice-covered, preventing deep convection and thus maintaining a weak AMOC. Deep water formation in the NorS ensures that the AMOC is not as weak as in the OFF state. In phase B, sea ice retreats in the LabS, allowing the ocean to release heat to the atmosphere and consequently deep convection to be activated. The salt-advection feedback kicks in, supplying warm and salty water to the LabS, enhancing sea ice retreat and convection up to the AMOC maximum in phase C. Then, however, the salinity and convection in the LabS start to decrease again, along with sea ice expansion. One possible reason for this reversal could be the strong freshwater influx from the Arctic Ocean through the Denmark Strait, which peaks in phase C and reaches to the LabS. The freshwater flow is concentrated in the upper ocean, implying that it can disrupt convection by freshening the upper water column. Another explanation could involve surface fluxes of heat and freshwater. In phase D, sea ice rapidly expands to cover the entire LabS and convection shuts down, causing the AMOC decline (Figs. 10d-g).
@@
Fully deciphering the oscillation mechanism goes beyond the scope of this study. Nonetheless, we can identify multiple competing processes that could produce cyclic behaviour: a competition between sea ice and convection in the LabS, a competition between salt advection by the AMOC and freshwater advection from the Arctic Ocean, as well as a competition of deep water formation sites between the LabS and NorS. The latter could explain the anti-phase pattern observed in precipitation between Greenland and the United Kingdom (Fig. 9g) as well as in surface air temperatures between the Greenland-Iceland-Norwegian (GIN) Seas and the rest of the high northern latitudes (Fig. 10c). During phase A, at the AMOC minimum, air temperatures are warmer over the GIN Seas than during phase C, in contrast with the surrounding areas, especially the LabS. We hypothesise that this is because the NorS is the only deep convection zone in the North Atlantic during phase A, directing the meridional heat transport to that region and reducing cold inflows from the Arctic Ocean.
@@
Exploring the dynamics on the edge state thus reveals distinct modes of climate variability that are absent in the ON and OFF attractors – but become highly relevant near criticality, as we show below. The edge tracking method allowed to capture the centennial climate oscillations even though they are asymptotically unstable.